OpenStudy (anonymous):

What is the missing number? Fifteen twenty-fifths is equal to three over what number? 3 5 twenty twenty-five

OpenStudy (anonymous):

I wanna say 5

OpenStudy (tanya123):

\(\displaystyle \frac{15}{25}=\frac{3}{x}\) Cross multiply and divide. answer 5